OpenStudy (anonymous):

When simplifying the expression 2/x + 2/x - 1 + 2/x + 2. The denominator is x(x-1)(x+2), what is the numerator?

OpenStudy (anonymous):

@amistre64

OpenStudy (espex):

How do you get the same denominator between two fractions?

OpenStudy (anonymous):

These are my choices 6x2 + 8x + 4 6x2 + 4x – 4 6x2 + 9x + 4 6x2 + 9x – 4

OpenStudy (espex):

To combine these, what is your first step? \[\frac{2}{x} + \frac{2}{x - 1} + \frac{2}{x + 2}\]

OpenStudy (anonymous):

@satellite73 walkthrough please

OpenStudy (yrelhan4):

 Its more or less like taking lcm of the denominators and solving..

OpenStudy (anonymous):

Walkthrough please?

OpenStudy (yrelhan4):

can you tell me how ill you do this.. 1/2 + 1/3 + 1/4 ?

OpenStudy (anonymous):

LCD so 12

OpenStudy (yrelhan4):

write it as i did.. in the drawing.

OpenStudy (anonymous):

I'm confused by your drawing from my understanding the denominator was already in the problem. It just didn't change... Clearly I messed up lol.